OpenStudy (anonymous):

Which of the following describes the rigid transformation in the function shown below? y+5= -2(x-1)^2 A. The graph is shifted 5 units down. B. The graph is shifted 1 unit left. C. The graph is reflected across the x-axis. D. The graph is stretched vertically by a factor of 2. ***idk please explain? thanks ;)

OpenStudy (anonymous):

do u understand this?

jimthompson5910 (jim_thompson5910):

i do, but more than one answer applies...so could there be more to the instructions?

OpenStudy (anonymous):

no :(

jimthompson5910 (jim_thompson5910):

well the +5 on the left side gets subtracted from both sides to get y = -2(x-1)^2 - 5 so that means choice A is the answer BUT the negative on -2 (outside the parenthesis) implies that the graph is reflected over the x-axis, so choice C is also the answer

OpenStudy (anonymous):

wait what??? so they are BOTH the answer???

jimthompson5910 (jim_thompson5910):

yeah that's exactly what I was thinking...they both can't be the answer

OpenStudy (anonymous):

which would be the most logical answer though? Answer A?

OpenStudy (anonymous):

well, last time it was reflected when it was like this kinda: y= -3x^2+1

OpenStudy (anonymous):

and this one would become -2x+2 so that would reflect also?

OpenStudy (anonymous):

and k=-5 so it is shifted 5 units down.. so idk? they both seem correct... but i can only choose one?

jimthompson5910 (jim_thompson5910):

yeah that's why I'm asking if there's more

OpenStudy (anonymous):

no that's all there is to the question

jimthompson5910 (jim_thompson5910):

hmm it's like they're doing this to mess with you lol

OpenStudy (anonymous):

does it mean anything when it says RIGID TRANSFORMATION?

OpenStudy (anonymous):

hahaa ikr?

jimthompson5910 (jim_thompson5910):

well they're all rigid transformations

OpenStudy (anonymous):

oh okay, so which one seems more logical to u tho?? i have to choose one now :/ and if it's wrong i can talk to my teacher :/

jimthompson5910 (jim_thompson5910):

oh wait, scaling isn't a rigid transformation

jimthompson5910 (jim_thompson5910):

so we can at least eliminate D

jimthompson5910 (jim_thompson5910):

it's either A or C

jimthompson5910 (jim_thompson5910):

B is incorrect because it's going the wrong way

OpenStudy (anonymous):

okay, which seems like the definite definite answer (and yes i meant to say definite twice teehehehe) :P

OpenStudy (anonymous):

? sorry that was a question haha

jimthompson5910 (jim_thompson5910):

hmm i'll go with C...it seems like we have to manipulate the equation a bit to get A

OpenStudy (anonymous):

okay =) ill just cross my fingers that C is the answer haha :) and i can always explain to my teacher later :)

OpenStudy (anonymous):

hahaa thanks for the help!! :D hope C is right!! :)

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (anonymous):

I'm not seeing the reflection across the x - axis, I think the answer is A.

OpenStudy (anonymous):

reflection across the x - axis implies that a given x value would result in two y values one above and one below the x - axis, but this is not the case

OpenStudy (anonymous):

what do you mean? can u pls explain some more?

OpenStudy (anonymous):

@tjones89 :)

OpenStudy (anonymous):

for a function to reflect about an axis, you should be able to plug in two different x values and get the same value for y, for example, y = x^2 if I plug in a -1 or a 1 I get the same y value

jimthompson5910 (jim_thompson5910):

you're thinking of a reflection across the y axis (not x axis)

OpenStudy (anonymous):

right but the same idea applies, if I want reflection about the x axis, I need to be able to use one x value to get two y values, one above and one below. This function does not have that property

jimthompson5910 (jim_thompson5910):

no function has that property...i think they want to know what transformations are done to the parent function y = x^2 to get y+5= -2(x-1)^2

OpenStudy (anonymous):

wait so my answer is C then still?

OpenStudy (anonymous):

but a reflection about the x axis implies that I should be able to fold the graph at the x axis and have the top and bottom parts line up, graphing this function shows that this is not the case

OpenStudy (anonymous):

 like this